# Hyperplane arrangements and Vinberg’s $\theta$-groups

### v1

@@ 2.2. Cartan subspaces and preliminary results
    We note that the order of $\theta$ does not play any role in this result, which is, in fact, true also in the case $m=1$. However, if $\theta$ is an inner automorphism of order $m\geq 2$, then there exists a lower bound on $W$ that depends on $m$. In fact, if $\theta$ is inner, there exists a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$ that is contained in $\mathfrak{g}_{0}$, see [32, §3.6]. Hence $\theta=\operatorname{Ad}_{g}$ for some $g\in Z_{G}(\mathfrak{h})=H\subset G_{0}$, where $H$ is the Cartan subgroup corresponding to $\mathfrak{h}$, and the class of $g$ in $W=N_{G_{0}}(\mathfrak{c})/Z_{G_{0}}(\mathfrak{c})$ has order $m$. Thus $W$ includes a subgroup isomorphic to $\mathbb{Z}/m\mathbb{Z}$.
@@ 3. Proof of the main result
This implies $G^{x}=Z_{G}(\mathfrak{c})$, since both centralizers are connected due to [30, Corollary 3.11]. In particular $N_{G_{0}}(\mathfrak{c})\cap G^{x}\subset Z_{G_{0}}(\mathfrak{c})$, whence the stabilizer of $x$ in $W$ is the trivial group. By Steinberg’s fixed point theorem for stabilizers in reflection groups [29, Theorem 1.5], this happens if and only if $x\notin\Pi_{s}$ for all $s\in\mathcal{R}(W)$.
@@
Choose $p\in\mathfrak{c}$ s.t. $\mathfrak{c}=\Pi_{\sigma}\oplus\mathbb{C}p$. By Claim I, we have $p\in\mathfrak{g}^{x}_{1}=\mathfrak{z}(\mathfrak{g}^{x})_{1}\oplus\mathfrak{m}_{1}=\Pi_{\sigma}\oplus\mathfrak{m}_{1}$ and thus we may write $p=z+p_{\sigma}$ with $z\in\Pi_{\sigma}$ and $p_{\sigma}\in\mathfrak{m}_{1}$. Because $q,z\in\mathfrak{c}$, then also $p_{\sigma}\in\mathfrak{c}$. From $0\neq\sigma(p)=\sigma(p_{\sigma})$ we conclude that $p_{\sigma}\neq 0$ and $\mathfrak{c}=\Pi_{\sigma}\oplus\mathbb{C}p_{\sigma}$, proving the claim.
@@
Choose $p_{\sigma}\in\mathfrak{m}$ s.t. $\mathfrak{c}=\Pi_{\sigma}\oplus\mathbb{C}p_{\sigma}$ as in Claim II. Clearly $\mathbb{C}p_{\sigma}\subset\mathfrak{m}_{1}$ consists of commuting semisimple elements, and we will now verify the two requirements as in [34, Proposition 6] for being a Cartan subspace.
The centralizer $\mathfrak{m}^{p_{\sigma}}$ is a Levi subalgebra of $\mathfrak{m}$ which is $\theta$-stable and admits the following decomposition
$\mathfrak{m}^{p_{\sigma}}=(\mathfrak{h}\cap\mathfrak{m})\oplus\bigoplus_{\alpha\in\Phi_{x},\alpha(p_{\sigma})=0}\mathfrak{g}_{\alpha}=(\mathfrak{h}\cap\mathfrak{m})\oplus\bigoplus_{\alpha\in\Phi,\alpha\circ\rho=0}\mathfrak{g}_{\alpha},$
with $\mathfrak{h}\cap\mathfrak{m}$ a $\theta$-stable Cartan subalgebra of $\mathfrak{m}^{p_{\sigma}}$. Then:
  1. (i)
    $\mathbb{C}p_{\sigma}\subset\mathfrak{z}(\mathfrak{m}^{p_{\sigma}})_{1}\subset(\mathfrak{h}\cap\mathfrak{m})_{1}=\mathfrak{c}\cap\mathfrak{m}=\mathbb{C}p_{\sigma}$, so that $\mathfrak{z}(\mathfrak{m}^{p_{\sigma}})_{1}=\mathbb{C}p_{\sigma}$;
  2. (ii)
    $\mathfrak{m}^{p_{\sigma}},\mathfrak{m}^{p_{\sigma}}]\subset[(\mathfrak{g}^{x})^{p_{\sigma}},(\mathfrak{g}^{x})^{p_{\sigma}}]=[\mathfrak{g}^{\mathfrak{c}},\mathfrak{g}^{\mathfrak{c}}]$, a periodically graded semisimple Lie algebra of zero rank (since $\mathfrak{c}$ is a Cartan subspace of $\mathfrak{g}$). Therefore $\mathfrak{m}^{p_{\sigma}},\mathfrak{m}^{p_{\sigma}}]$ has zero rank too.
The two requirements are thus established and so is Claim III.
@@ 4.3. Inner automorphisms of $\mathfrak{sl}_{n+1}(\mathbb{C})$
By the results of [34, §7-§8] and [15, §8], any inner automorphism $\theta:\mathfrak{g}\to\mathfrak{g}$ of $\mathfrak{g}=\mathfrak{sl}_{n+1}(\mathbb{C})$ of positive rank is described, up to conjugation, by a crossed Kac diagram. This is the affine Dynkin diagram of $\mathfrak{g}$, where each node has an additional label $1$, if the node is crossed, or 0, otherwise. The order $m$ of $\theta$ is the number of crossed nodes, so the Kac diagrams we are considering have always at least one cross and we may assume w.l.o.g. that this is placed on the lowest root $\alpha_{0}$. In summary, we have an ordered partition $(k_{1},k_{2},\ldots,k_{N})$ of $n+1$, where $k_{1}$ is the number of consecutive crossed nodes counting from $\alpha_{0}$, $k_{2}$ is the number of consecutive uncrossed nodes counting from the simple root $\alpha_{k_{1}}$, etc. Thanks to [34, Proposition 17], $\mathfrak{g}_{0},\mathfrak{g}_{0}]\cong\mathfrak{sl}_{k_{2}+1}(\mathbb{C})\oplus\mathfrak{sl}_{k_{4}+1}(\mathbb{C})\oplus\cdots$ and $\mathfrak{g}_{1}$ is the sum of $m$ irreducible representations of $\mathfrak{g}_{0}$ (one for each crossed node among the following representations $\mathbb{C}$, $\mathbb{C}^{k_{2j}+1}$, $(\mathbb{C}^{k_{2j}+1})^{*}$, or $\mathbb{C}^{k_{2j}+1}\otimes(\mathbb{C}^{k_{2(j+1)}+1})^{*}$).
@@
It can be shown that the rank $r$ is equal to the minimum of the eigenvalue multiplicities thanks to [34, §7]. This then suggests to perform an additional conjugation, varying $x$ in $\mathfrak{g}$ in such a way that $\theta$ is determined by the block diagonal matrix
@@
The triple $\{\mathfrak{s}:=m\,\mathfrak{sl}_{2}(\mathbb{C}),\theta,m\}$ fits thus into the discussion carried out in [34, §1.2, §2.2], with $\{\mathfrak{sl}_{2}(\mathbb{C}),\operatorname{id},1\}$ as the associated simple component. It follows (see, e.g., [34, pag. 479]) that the little Weyl group of the triple, considered as a linear group, is isomorphic to the Weyl group of the simple component. The latter is $\mathbb{Z}_{2}$ and the sought lift of the reflection is thus $\exp(\pi J)$ with $J:=\frac{1}{2}\big{(}e_{+}^{(0)}-e_{-}^{(0)}\big{)}\in\mathfrak{s}_{0}$, arguing as in §4.2. This concludes the discussion of case (2) and the proof of Theorem 4.4.
@@ Acknowledgments
The authors would like to thank Giovanna Carnovale, Willem de Graaf, Oksana Yakimova for helpful and stimulating discussions. The first author’s research work was funded by FSU Jena. The second author acknowledges the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006. This article/publication was also supported by the “National Group for Algebraic and Geometric Structures, and their Applications” GNSAGA-INdAM (Italy) and it is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology), https://www.cost.eu.

### v2

@@ 2.2. Cartan subspaces and preliminary results
    We note that the order of $\theta$ does not play any role in this result, which is, in fact, true also in the case $m=1$. However, if $\theta$ is an inner automorphism of order $m\geq 2$, then there exists a lower bound on $W$ that depends on $m$. In fact, if $\theta$ is inner, there exists a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$ that is contained in $\mathfrak{g}_{0}$, see [31, §3.6]. Hence $\theta=\operatorname{Ad}_{g}$ for some $g\in Z_{G}(\mathfrak{h})=H\subset G_{0}$, where $H$ is the Cartan subgroup corresponding to $\mathfrak{h}$, and the class of $g$ in $W=N_{G_{0}}(\mathfrak{c})/Z_{G_{0}}(\mathfrak{c})$ has order $m$. Thus $W$ includes a subgroup isomorphic to $\mathbb{Z}/m\mathbb{Z}$. This bound has already been observed in [27, Lemma 23].
@@ 3. Proof of the main result
This implies $G^{x}=Z_{G}(\mathfrak{c})$, since both centralizers are connected due to [29, Corollary 3.11]. In particular $N_{G_{0}}(\mathfrak{c})\cap G^{x}\subset Z_{G_{0}}(\mathfrak{c})$, whence the stabilizer of $x$ in $W$ is the trivial group, and $x\notin\Pi_{s}$ for all $s\in\mathcal{R}(W)$.
@@
Choose $p\in\mathfrak{c}$ s.t. $\mathfrak{c}=\Pi_{\sigma}\oplus\mathbb{C}p$. By Claim I, we have $p\in\mathfrak{g}^{x}_{1}=\mathfrak{z}(\mathfrak{g}^{x})_{1}\oplus\mathfrak{m}_{1}=\Pi_{\sigma}\oplus\mathfrak{m}_{1}$ and thus we may write $p=z+p_{\sigma}$ with $z\in\Pi_{\sigma}$ and $p_{\sigma}\in\mathfrak{m}_{1}$. Because $p,z\in\mathfrak{c}$, then also $p_{\sigma}\in\mathfrak{c}$. From $0\neq\sigma(p)=\sigma(p_{\sigma})$ we conclude that $p_{\sigma}\neq 0$ and $\mathfrak{c}=\Pi_{\sigma}\oplus\mathbb{C}p_{\sigma}$, proving the claim.
@@
Choose $p_{\sigma}\in\mathfrak{m}$ s.t. $\mathfrak{c}=\Pi_{\sigma}\oplus\mathbb{C}p_{\sigma}$ as in Claim II. Clearly $\mathbb{C}p_{\sigma}\subset\mathfrak{m}_{1}$ consists of commuting semisimple elements, so it lies in a Cartan subspace $\mathfrak{c}_{\mathfrak{m}}$ for $\mathfrak{m}$. Since $\mathfrak{m},\Pi_{\sigma}]=0$, the subspace $\mathfrak{c}_{\mathfrak{m}}\oplus\Pi_{\sigma}$ is abelian and consists of semisimple elements. Hence $\dim(\mathfrak{c}_{\mathfrak{m}}\oplus\Pi_{\sigma})\geq\dim\Pi_{\sigma}+1$, which agrees with the rank of $\{\mathfrak{g},\theta\}$, and $\mathfrak{c}_{\mathfrak{m}}\oplus\Pi_{\sigma}$ is a Cartan subspace of $\{\mathfrak{g},\theta\}$ by maximality. In summary $\dim\mathfrak{c}_{\mathfrak{m}}=1$ and $\mathfrak{c}_{\mathfrak{m}}=\mathbb{C}p_{\alpha}$, proving Claim III.
@@ 4.3. Inner automorphisms of $\mathfrak{sl}_{n+1}(\mathbb{C})$
By the results of [33, §7-§8] and [15, §8], any inner automorphism $\theta:\mathfrak{g}\to\mathfrak{g}$ of $\mathfrak{g}=\mathfrak{sl}_{n+1}(\mathbb{C})$ of positive rank is described, up to conjugation, by a crossed Kac diagram. This is the affine Dynkin diagram of $\mathfrak{g}$, where each node has an additional label $1$, if the node is crossed, or 0, otherwise. The order $m$ of $\theta$ is the number of crossed nodes, so the Kac diagrams we are considering have always at least one cross and we may assume w.l.o.g. that this is placed on the lowest root $\alpha_{0}$. In summary, we have an ordered partition $(k_{1},k_{2},\ldots,k_{N})$ of $n+1$, where $k_{1}$ is the number of consecutive crossed nodes counting from $\alpha_{0}$, $k_{2}$ is the number of consecutive uncrossed nodes counting from the simple root $\alpha_{k_{1}}$, etc. Thanks to [33, Proposition 17], $\mathfrak{g}_{0},\mathfrak{g}_{0}]\cong\mathfrak{sl}_{k_{2}+1}(\mathbb{C})\oplus\mathfrak{sl}_{k_{4}+1}(\mathbb{C})\oplus\cdots$ and $\mathfrak{g}_{1}$ is the sum of $m$ irreducible representations of $\mathfrak{g}_{0}$ (one for each crossed node among the following representations $\mathbb{C}$, $\mathbb{C}^{k_{2j}+1}$, $(\mathbb{C}^{k_{2j}+1})^{*}$, or $\mathbb{C}^{k_{2j}+1}\otimes(\mathbb{C}^{k_{2(j+1)}+1})^{*}$).
@@
It can be shown that the rank $r$ is equal to the minimum of the eigenvalue multiplicities thanks to [33, §7]. This then suggests to perform an additional conjugation, varying $x$ in $\mathfrak{g}$ in such a way that $\theta$ is determined by the block diagonal matrix
@@
The triple $\{\mathfrak{s}:=m\,\mathfrak{sl}_{2}(\mathbb{C}),\theta,m\}$ fits thus into the discussion carried out in [33, §1.2, §2.2], with $\{\mathfrak{sl}_{2}(\mathbb{C}),\operatorname{id},1\}$ as the associated simple component. It follows (see, e.g., [33, pag. 479]) that the little Weyl group of the triple, considered as a linear group, is isomorphic to the Weyl group of the simple component. The latter is $\mathbb{Z}_{2}$ and the sought lift of the reflection is thus $\exp(\pi J)$ with $J:=\frac{1}{2}\big(e_{+}^{(0)}-e_{-}^{(0)}\big)\in\mathfrak{s}_{0}$, arguing as in §4.2. This concludes the discussion of case (2) and the proof of Theorem 4.4.
@@ Acknowledgments
The authors would like to thank Giovanna Carnovale, Willem de Graaf, Francesco Esposito, and Oksana Yakimova for helpful and stimulating discussions. The authors also thank the anonymous referee for carefully reading the manuscript and providing helpful comments that improved the paper. The first author’s research work was funded by FSU Jena. The second author acknowledges the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006. This article/publication was also supported by the “National Group for Algebraic and Geometric Structures, and their Applications” GNSAGA-INdAM (Italy) and it is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology), https://www.cost.eu.